# A route to quantum computing through the theory of quantum graphs

### v2

@@ Abstract
Based on our previous works, and in order to relate them with the theory of quantum graphs and the quantum computing principles, we once again try to introduce some newly developed technical structures just by relying on our toy example, i.e. the coordinate ring of $n\times n$ quantum matrix algebra $M_{q}(n)$, and the associated directed locally finite graphs $\mathcal{G}(\Pi_{n})$, and the Cuntz-Krieger $C^{*}$-graph algebras. Meaningly, we introduce a $(4i-6)$-qubit quantum system by using the Cuntz-Krieger $\mathcal{G}(\Pi_{i})$-families associated to the $64i-$ 6 distinct Hamiltonian paths of $\mathcal{G}(\Pi_{i})$, for $i\in\{2,\cdots,n\}$.
@@ 4 $C^{*}$-Graph Algebras, a rout to quantum ($C^{*}$-algebraic) graphs
$S_{g}:=\sum_{n=1}^{\infty}E_{6n-4,3n-1},S_{i}:=\sum_{n=1}^{\infty}E_{6n-1,3n},S_{j}:=\sum_{n=1}^{\infty}E_{6n-3,3n-1}\}$ (4.2)
@@ 4.1 A route to quantum computing through the theory of quantum graphs
Then by employing the defining relations (4.2), it will not be too difficult to see that $S_{e}^{*}S_{e}+S_{i}^{*}S_{i}+S_{g}^{*}S_{g}=I$, (and $S_{f}^{*}S_{f}+S_{j}^{*}S_{j}+S_{h}^{*}S_{h}=I$) are true statements within $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ respectively, and then by using the definition 3.3, we might be interested in defining a completely positive trace-preserving map $\Psi_{\mathcal{G}(\Pi_{2})}^{1}:M_{m}\to M_{m}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}:M_{m}\to M_{m}$); where $m$ will be specified later; in a way that we have $\Psi_{\mathcal{G}(\Pi_{2})}^{1}(\mathcal{X})=\sum_{m}S_{m}\mathcal{X}S_{m}^{*}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}(\mathcal{X})=\sum_{m}S_{m}\mathcal{X}S_{m}^{*}$) for all $\mathcal{X}\in M_{m}$, and we will call it our quantum channel. After that, for $\ell\in\{e,g,i\}$ (and $\ell\in\{f,h,j\}$), the generating partial (matrix) isometries $S_{\ell}$, will be called the Choi-Kraus operators, and the noncommutative (confusability) graph of $\Psi_{\mathcal{G}(\Pi_{2})}^{1}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}$) will be the operator system
@@
Postulate 4.9 (State Space Postulate).
@@
Here, our quantum system will be a combination of two subsystems $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$ with the correspondent states $\phi_{1}$ and $\phi_{2}$, and since in this case, the Hamiltonian paths are a combination of three edges, hence let $(\mathbb{C}^{2})^{\otimes 1}\cong\mathbb{C}^{2}$ be our state space $\phi$. Consider $\phi_{1}=\alpha_{1}|0\rangle+\alpha_{2}|1\rangle$ and $\phi_{2}=\beta_{1}|0\rangle+\beta_{2}|1\rangle$ such that we have $\sum_{i=1}^{2}\alpha_{i}^{2}=\sum_{i=1}^{2}\beta_{i}^{2}=1$. Not that the only possible options for $\alpha_{i}$s and $\beta_{i}$s are ±$\pm\frac{1}{\sqrt{2}}$.
@@
$\text{where}\ \sum_{j\in\{0,1\}^{2}}|\alpha_{j}|^{2}=1$
@@
It is easy to see that, in (4.16), the only possible options for $\alpha_{k}$s are ±$\pm\frac{1}{2}$, and as a result we have the following proposition.
@@
The proof of this proposition is almost a trivial conclusion of the Definition 4.18, because, in any case, $|\phi\rangle$, and $|\psi\rangle$, respectively for $\mathcal{G}(\Pi_{2})$ and $\mathcal{G}(\Pi_{3})$, will always be equal to the tensor product $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$, because in $\mathcal{G}(\Pi_{2})$ the amplitudes $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$ will be equal to ±$\pm\frac{1}{2}$, just the same as in $|\phi\rangle$, and in $\mathcal{G}(\Pi_{3})$ the amplitudes $|\psi_{1}\rangle\otimes|\psi_{2}\rangle$ will be equal to ±$\pm\frac{1}{2^{3}}$, just the same as in $|\psi\rangle$
@@
Note that in order to prove or disprove this claim, the person first needs to prove or disprove Claim 2.7., from [7].

### v3

@@ Abstract
Based on our previous works, and in order to relate them with the theory of quantum graphs and the quantum computing principles, we once again try to introduce some newly developed technical structures just by relying on our toy example, the coordinate ring of $n\times n$ quantum matrix algebra $M_{q}(n)$, and the associated directed locally finite graphs $\mathcal{G}(\Pi_{n})$, and the Cuntz-Krieger $C^{*}$-graph algebras. Meaningly, we introduce a $(4i-6)$-qubit quantum system by using the Cuntz-Krieger $\mathcal{G}(\Pi_{i})$-families associated to the $64i-$ 6 distinct Hamiltonian paths of $\mathcal{G}(\Pi_{i})$, for $i\in\{2,\cdots,n\}$. We also will present a proof of a claim raised in our previous paper concerning the graph $C^{*}$-algebra structure and the associated Cuntz-Krieger $\mathcal{G}(\Pi_{n})$-families.
@@ 4 $C^{*}$-Graph Algebras, a rout to quantum ($C^{*}$-algebraic) graphs
$S_{g}:=\sum_{n=1}^{\infty}E_{6n-4,3n-1},S_{i}:=\sum_{n=1}^{\infty}E_{6n-1,3n},S_{j}:=\sum_{n=1}^{\infty}E_{6n-3,3n-1}\},$ (4.2)
@@ 4.1 A route to quantum computing through the theory of quantum graphs
Then by employing the defining relations (4.21),
it will not be too difficult to see that $S_{e}^{*}S_{e}+S_{i}^{*}S_{i}+S_{g}^{*}S_{g}=I$, (and $S_{f}^{*}S_{f}+S_{j}^{*}S_{j}+S_{h}^{*}S_{h}=I$) are true statements within $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ respectively, and then by using the definition 3.3, we might be interested in defining a completely positive trace-preserving map $\Psi_{\mathcal{G}(\Pi_{2})}^{1}:M_{m}\to M_{m}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}:M_{m}\to M_{m}$); where $m$ will be specified later; in a way that we have $\Psi_{\mathcal{G}(\Pi_{2})}^{1}(\mathcal{X})=\sum_{m}S_{m}\mathcal{X}S_{m}^{*}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}(\mathcal{X})=\sum_{m}S_{m}\mathcal{X}S_{m}^{*}$) for all $\mathcal{X}\in M_{m}$, and we will call it our quantum channel. After that, for $\ell\in\{e,g,i\}$ (and $\ell\in\{f,h,j\}$), the generating partial (matrix) isometries $S_{\ell}$, will be called the Choi-Kraus operators, and the noncommutative (confusability) graph of $\Psi_{\mathcal{G}(\Pi_{2})}^{1}$ (and $\Psi_{\mathcal{G}(\Pi_{2})}^{2}$) will be the operator system
@@
Let us try to have some estimations for the quantum channels from the Lemma 4.5.
For any $\mathcal{X}\in M_{4}$, we have
$\Psi_{\mathcal{G}(\Pi_{2})}^{1}(\mathcal{X})$ $=\sum\limits_{m\in\{e,i,g\}}S_{m}\mathcal{X}S_{m}^{*}$
$=S_{e}\mathcal{X}S_{e}^{*}+S_{i}\mathcal{X}S_{i}^{*}+S_{g}\mathcal{X}S_{g}^{*}$
$=\sum\limits_{n=1}^{\infty}E_{6n,3n-2}\mathcal{X}\sum\limits_{n=1}^{\infty}E_{3n-2,6n}+\sum\limits_{n=1}^{\infty}E_{6n-1,3n}\mathcal{X}\sum\limits_{n=1}^{\infty}E_{3n,6n-1}$
$\hskip 8.5359pt+\sum\limits_{n=1}^{\infty}E_{6n-4,3n-1}\mathcal{X}\sum\limits_{n=1}^{\infty}E_{3n-1,6n-4}$
$=x_{22}\sum\limits_{n=1}^{\infty}E_{3n-1,6n-4}$
$=x_{22}(E_{2,2}+E_{5,8}+E_{8,14}+\cdots).$
And for $\Psi_{\mathcal{G}(\Pi_{2})}^{2}$ we have
$\Psi_{\mathcal{G}(\Pi_{2})}^{2}(\mathcal{X})$ $=\sum\limits_{m\in\{f,j,h\}}S_{m}\mathcal{X}S_{m}^{*}$
$=S_{f}\mathcal{X}S_{f}^{*}+S_{j}\mathcal{X}S_{j}^{*}+S_{h}\mathcal{X}S_{h}^{*}$
$=\sum\limits_{n=1}^{\infty}E_{6n-4,3n-2}\mathcal{X}\sum\limits_{n=1}^{\infty}E_{3n-2,6n-4}+\sum\limits_{n=1}^{\infty}E_{6n-3,3n-1}\mathcal{X}\sum\limits_{n=1}^{\infty}E_{3n-1,6n-3}$
$\hskip 8.5359pt+\sum\limits_{n=1}^{\infty}E_{6n-3,3n}\mathcal{X}\sum\limits_{n=1}^{\infty}E_{3n,6n-3}$
$=x_{22}\sum\limits_{n=1}^{\infty}E_{3n-2,6n-4}+x_{32}\sum\limits_{n=1}^{\infty}E_{3n-1,6n-3}+x_{33}\sum\limits_{n=1}^{\infty}E_{3n,6n-3}$
$=x_{22}(E_{1,2}+E_{4,8}+E_{7,14}+\cdots)+x_{32}(E_{2,3}+E_{5,9}+E_{8,15}+\cdots)$
$\hskip 8.5359pt+x_{33}(E_{3,3}+E_{6,9}+E_{9,15}+\cdots).$
Question 4.6.
So, now the question is that, what do these estimations tell us?
@@
Postulate 4.10 (State Space Postulate).
@@
Here, our quantum system will be a combination of two subsystems $\mathbb{E}_{1}$ and $\mathbb{E}_{2}$ with the correspondent states $\phi_{1}$ and $\phi_{2}$, and since in this case, the Hamiltonian paths are a combination of three edges, hence let $(\mathbb{C}^{2})^{\otimes 1}\cong\mathbb{C}^{2}$ be our state space $\phi$. Consider $\phi_{1}=\alpha_{1}|0\rangle+\alpha_{2}|1\rangle$ and $\phi_{2}=\beta_{1}|0\rangle+\beta_{2}|1\rangle$ such that we have $\sum_{i=1}^{2}\alpha_{i}^{2}=\sum_{i=1}^{2}\beta_{i}^{2}=1$. Not that the only possible options for $\alpha_{i}$s and $\beta_{i}$s are ±$\pm\frac{1}{\sqrt{2}}$ and ±$\pm\frac{i}{\sqrt{2}}$.
@@
$\text{where}\ \sum_{j\in\{0,1\}^{2}}|\alpha_{j}|^{2}=1,$
@@
It is easy to see that, in (4.17), the only possible options for $\alpha_{k}$s are ±$\pm\frac{1}{2}$ and ±$\pm\frac{i}{2}$, and as a result we have the following proposition.
@@
The proof of this proposition is almost a trivial conclusion of the Definition 4.19, because, in any case, $|\phi\rangle$, and $|\psi\rangle$, respectively for $\mathcal{G}(\Pi_{2})$ and $\mathcal{G}(\Pi_{3})$, will always be equal to the tensor product $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$, because in $\mathcal{G}(\Pi_{2})$ the amplitudes $|\phi_{1}\rangle\otimes|\phi_{2}\rangle$ will be equal to ±$\pm\frac{1}{2}$ and ±$\pm\frac{i}{2}$, just the same as in $|\phi\rangle$, and in $\mathcal{G}(\Pi_{3})$ the amplitudes $|\psi_{1}\rangle\otimes|\psi_{2}\rangle$ will be equal to ±$\pm\frac{1}{2^{3}}$ and ±$\pm\frac{i}{2^{3}}$, just the same as in $|\psi\rangle$
@@
Note that in order to prove or disprove this claim, the interested person first needs to prove or disprove Claim 2.7., from [7], on which we will try to present a sketch of its proof here in this paper as the following Theorem.
Theorem 4.21.
For $\mathcal{G}(\Pi_{n})=(\mathcal{G}^{0},\mathcal{G}^{1})$ the associated directed locally finite graphs with $\mathbb{K}[M_{q}(n)]$, and $\Pi_{n}$ the associated adjacency matrices, and $\mathcal{H}:=\ell^{2}(\mathbb{N})$ the underlying infinite dimensional Hilbert space. The claim is that the set
$S=\{S_{i}:=\sum_{j=1}^{\infty}\prescript{i}{}{E}_{\mathcal{E}j-A,(n^{2}-1)j-D}\mid\ \text{for fixed}\ 1\leq i\leq\frac{(n^{3}+n^{2})(n-1)}{2}\},$
is a Cuntz-Krieger $\mathcal{G}(\Pi_{n})$-family for $D\in\{0,\cdots,n^{2}-2\}$, and $\mathcal{E}$ depends on the degree of the exit edges to the vertex $e_{hk}$, where $i$ is considered as an exit edge, i.e. if $\overset{\rightarrow}{\,\mathsf{deg}\,}_{hk}=2$, then we will have $\mathcal{E}=2(n^{2}-1)$, and if it is 3, then we will have $\mathcal{E}=3(n^{2}-1)$, and so on, and $A\in\{0,\cdots,\overset{\rightarrow}{\,\mathsf{deg}\,}_{hk}\times(n^{2}-1)\}$, and gives us a graph $C^{*}$-algebra structure $\mathcal{C}^{*}(\Pi_{n})$.
Proof.
In order to prove, we will proceed by induction. But before proceeding, let us clarify the almost approximate values of the parameters $D$. Starting with Hamiltonian paths, we know that they start from the only source vertex $e_{1,1}$ and terminate at the only sink vertex $e_{n,n}$. Hence depending on the first connection, we might have each of $(n^{2}-1)i-D$, for $D\in\{0,\cdots,n^{2}-2\}$. For example, if the first connection is $e_{1,1}\rightarrow e_{1,2}$, then we have $D=0$, or if it is $e_{1,1}\rightarrow e_{2,1}$, then we have $D=1$, and etc.
In [7], we already have proved that the assertion is true for $n=2$. Bellow, we will prove that it also is true for $n=3$.
Exactly in a same approach as in [7, Proposition 3.1], we note that there might be finite or infinite dimensional sets of projections. But in order to skip the complication, and to shorten the proof, we skip proving that for some certain vertices $e_{i,j}$, we should have $\dim(P_{e_{i,j}})>\infty$, and based on an already proven fact, all the other projections also should be infinite dimensional. So, let us try to construct an appropriate CK-$\mathcal{G}(\Pi_{3})$ family satisfying the relations $S_{e}^{*}S_{e}=P_{r(e)}$, for all edges $e\in\mathcal{G}(\Pi_{3})^{1}$, and $P_{e_{ij}}=\sum\limits_{s(e)=e_{i,j}}S_{e}S_{e}^{*}$ for the case when $e_{i,j}\in\mathcal{G}(\Pi_{3})^{0}$ is not a sink.
$e_{21}$$e_{22}$$e_{23}$$e_{31}$$e_{11}$$e_{33}$$e_{12}$$e_{13}$$e_{32}$$j_{5}$
$g_{6}=f_{6}$
$e_{3}=f_{3}$
$g_{1}=h_{1}$
$i_{8}=e_{8}=j_{8}=h_{8}$
[…]
$e_{1}=f_{1}$
[…]
$g_{4}=h_{4}=j_{2}$
$e_{4}=f_{4}=i_{2}$
[…]
$e_{5}=f_{5}$
$e_{7}=h_{7}=i_{7}=j_{7}$
$g_{7}=f_{7}$
[…]
Figure 2: Directed locally connected graph $\mathcal{G}(\Pi_{3})$
We note that $\mathcal{G}(\Pi_{3})$ consists of $6$ Hamiltonian paths listed below:
$\mathcal{H}_{1}:e_{11}\overset{e_{1}}{\longrightarrow}e_{12}\overset{e_{2}}{\longrightarrow}e_{21}\overset{e_{3}}{\longrightarrow}e_{31}\overset{e_{4}}{\longrightarrow}e_{13}\overset{e_{5}}{\longrightarrow}e_{22}\overset{e_{6}}{\longrightarrow}e_{32}\overset{e_{7}}{\longrightarrow}e_{23}\overset{e_{8}}{\longrightarrow}e_{33},$
$\mathcal{H}_{2}:e_{11}\overset{f_{1}}{\longrightarrow}e_{12}\overset{f_{2}}{\longrightarrow}e_{21}\overset{f_{3}}{\longrightarrow}e_{31}\overset{f_{4}}{\longrightarrow}e_{13}\overset{f_{5}}{\longrightarrow}e_{22}\overset{f_{6}}{\longrightarrow}e_{23}\overset{f_{7}}{\longrightarrow}e_{32}\overset{f_{8}}{\longrightarrow}e_{33},$
$\mathcal{H}_{3}:e_{11}\overset{g_{1}}{\longrightarrow}e_{21}\overset{g_{2}}{\longrightarrow}e_{12}\overset{g_{3}}{\longrightarrow}e_{13}\overset{g_{4}}{\longrightarrow}e_{31}\overset{g_{5}}{\longrightarrow}e_{22}\overset{g_{6}}{\longrightarrow}e_{23}\overset{g_{7}}{\longrightarrow}e_{32}\overset{g_{8}}{\longrightarrow}e_{33},$
$\mathcal{H}_{4}:e_{11}\overset{h_{1}}{\longrightarrow}e_{21}\overset{h_{2}}{\longrightarrow}e_{12}\overset{h_{3}}{\longrightarrow}e_{13}\overset{h_{4}}{\longrightarrow}e_{31}\overset{h_{5}}{\longrightarrow}e_{22}\overset{h_{6}}{\longrightarrow}e_{32}\overset{h_{7}}{\longrightarrow}e_{23}\overset{h_{8}}{\longrightarrow}e_{33},$
$\mathcal{H}_{5}:e_{11}\overset{i_{1}}{\longrightarrow}e_{31}\overset{i_{2}}{\longrightarrow}e_{13}\overset{i_{3}}{\longrightarrow}e_{21}\overset{i_{4}}{\longrightarrow}e_{12}\overset{i_{5}}{\longrightarrow}e_{22}\overset{i_{6}}{\longrightarrow}e_{32}\overset{i_{7}}{\longrightarrow}e_{23}\overset{i_{8}}{\longrightarrow}e_{33},$
$\mathcal{H}_{6}:e_{11}\overset{j_{1}}{\longrightarrow}e_{13}\overset{j_{2}}{\longrightarrow}e_{31}\overset{j_{3}}{\longrightarrow}e_{12}\overset{j_{4}}{\longrightarrow}e_{21}\overset{j_{5}}{\longrightarrow}e_{22}\overset{j_{6}}{\longrightarrow}e_{32}\overset{j_{7}}{\longrightarrow}e_{23}\overset{j_{8}}{\longrightarrow}e_{33}.$
Now the claim is that the desired set of partial isometries will be as follows:
$S_{e_{1}}=S_{f_{1}}=E_{32n-31,8n},S_{g_{1}}=S_{h_{1}}=E_{32n-23,8n-1},S_{i_{1}}=E_{32n-7,8n-2},$
$S_{j_{1}}=E_{32n-15,8n-3},$
$S_{g_{2}}=S_{h_{2}}=S_{i_{4}}=E_{40n-25,8n},S_{e_{2}}=S_{f_{2}}=S_{j_{4}}=E_{40n-32,8n-1},S_{j_{2}}=E_{48n-35,8n-2},$
$S_{i_{2}}=E_{48n-42,8n-3},$
$S_{j_{3}}=E_{48n-10,8n},S_{i_{3}}=E_{48n-27,8n-1},S_{e_{3}}=S_{f_{3}}=E_{40n-33,8n-2},$
$S_{g_{3}}=S_{h_{3}}=E_{40n-24,8n-3},$
$S_{i_{4}}=E_{40n-25,8n},S_{j_{4}}=E_{40n-32,8n-1},S_{g_{4}}=S_{h_{4}}=S_{j_{2}}=E_{48n-35,8n-2},$
$S_{e_{4}}=S_{f_{4}}=S_{i_{2}}=E_{48n-42,8n-3},$
$S_{e_{5}}=S_{f_{5}}=E_{48n-43,8n-4},S_{g_{5}}=S_{h_{5}}=E_{48n-34,8n-4},S_{i_{5}}=E_{40n-8,8n-4},$
$S_{j_{5}}=E_{40n-1,8n-4},$
$S_{e_{6}}=S_{h_{6}}=S_{i_{6}}=S_{j_{6}}=E_{32n-28,8n-5},S_{g_{6}}=S_{f_{6}}=E_{24n-14,8n-5},S_{i_{6}}=E_{32n-28,8n-5},$
$S_{j_{6}}=E_{32n-28,8n-5},$
$S_{e_{7}}=S_{h_{7}}=S_{i_{7}}=S_{j_{7}}=E_{24n-21,8n-6},S_{g_{7}}=S_{f_{7}}=E_{32n-20,8n-6},S_{i_{7}}=E_{24n-21,8n-6},$
$S_{j_{7}}=E_{24n-21,8n-6},$
$S_{e_{8}}=S_{h_{8}}=S_{i_{8}}=S_{j_{8}}=E_{24n-22,8n-7},S_{g_{8}}=S_{f_{8}}=E_{24n-13,8n-7},S_{i_{8}}=E_{24n-22,8n-7},$
$S_{j_{8}}=E_{24n-22,8n-7},$
and the rest of the isometries are apart from the set of Hamiltonian paths, and could be outlined as follows:
$S_{1}=E_{48n-2,8n-7},S_{14}=E_{48n-11,8n-7},S_{4}=E_{48n-29,8n-7},$
$S_{7}=E_{48n-19,8n-6},S_{12}=E_{40n-9,8n-6},$
$S_{2}=E_{48n-3,8n-5},S_{11}=E_{48n-26,8n-5},S_{13}=E_{40n-11,8n-5},$
$S_{3}=E_{24n-5,8n-3},S_{8}=E_{40n-17,8n-3},S_{9}=E_{32n-12,8n-3},$
$S_{5}=E_{24n-6,8n-2},S_{6}=E_{40n-16,8n-2},S_{5}=E_{32n-4,8n-2}.$
Now the goal is to prove that the above isometries satisfy in the relations of Cuntz-Krieger $\mathcal{G}(\Pi_{3})$-family, i.e. $S_{e}^{*}S_{e}=P_{r(e)}$, for all edges $e\in\mathcal{G}(\Pi_{3})^{1}$, and $P_{e_{ij}}=\sum\limits_{s(e)=e_{i,j}}S_{e}S_{e}^{*}$ for the case when $e_{i,j}\in\mathcal{G}(\Pi_{3})^{0}$ is not a sink, and this could be verified simply. Here we just put it over as follows.
$P_{e_{13}}=S_{3}^{*}S_{3}=S_{g_{3}}^{*}S_{g_{3}}=S_{j_{1}}^{*}S_{j_{1}}=S_{e_{4}}^{*}S_{e_{4}}=S_{9}^{*}S_{9}=S_{8}^{*}S_{8}=S_{e_{5}}S_{e_{5}}^{*}+S_{g_{4}}S_{g_{4}}^{*}+S_{i_{3}}S_{i_{3}}^{*}$
$\hskip 19.91684pt+S_{7}S_{7}^{*}+S_{14}S_{14}^{*}+S_{2}S_{2}^{*},$
$P_{e_{22}}=S_{e_{5}}^{*}S_{e_{5}}=S_{g_{5}}^{*}S_{g_{5}}=S_{j_{5}}^{*}S_{j_{5}}=S_{i_{5}}^{*}S_{i_{5}}=S_{e_{6}}S_{e_{6}}^{*}+S_{g_{6}}S_{g_{6}}^{*}+S_{9}S_{9}^{*}+S_{10}S_{10}^{*},$
$P_{e_{23}}=S_{e_{7}}^{*}S_{e_{7}}=S_{12}^{*}S_{12}=S_{g_{6}}^{*}S_{g_{6}}=S_{7}^{*}S_{7}=S_{4}^{*}S_{4}=S_{e_{8}}S_{e_{8}}^{*}+S_{g_{7}}S_{g_{7}}^{*}+S_{5}S_{5}^{*},$
$P_{e_{32}}=S_{2}^{*}S_{2}=S_{g_{7}}^{*}S_{g_{7}}=S_{e_{6}}^{*}S_{e_{6}}=S_{13}^{*}S_{13}=S_{11}^{*}S_{11}=S_{e_{7}}S_{e_{7}}^{*}+S_{g_{8}}S_{g_{8}}^{*}+S_{3}S_{3}^{*},$
$P_{e_{33}}=S_{1}^{*}S_{1}=S_{14}^{*}S_{14}=S_{e_{8}}^{*}S_{e_{8}}=S_{g_{8}}^{*}S_{g_{8}},$
$P_{e_{31}}=S_{i_{1}}^{*}S_{i_{1}}=S_{e_{3}}^{*}S_{e_{3}}=S_{10}^{*}S_{10}=S_{g_{4}}^{*}S_{g_{4}}=S_{5}^{*}S_{5}=S_{6}^{*}S_{6}=S_{e_{4}}S_{e_{4}}^{*}+S_{g_{5}}S_{g_{5}}^{*}+S_{11}S_{11}^{*}$
$\hskip 19.91684pt+S_{4}S_{4}^{*}+S_{j_{3}}S_{j_{3}}^{*}+S_{1}S_{1}^{*},$
$P_{e_{11}}=S_{e_{1}}^{*}S_{e_{1}}+S_{g_{1}}S_{g_{1}}^{*}+S_{i_{1}}S_{i_{1}}^{*}+S_{j_{1}}S_{j_{1}}^{*},$
$P_{e_{21}}=S_{g_{1}}^{*}S_{g_{1}}=S_{e_{2}}^{*}S_{e_{2}}=S_{i_{3}}^{*}S_{i_{3}}=S_{e_{3}}S_{e_{3}}^{*}+S_{g_{2}}S_{g_{2}}^{*}+S_{8}S_{8}^{*}+S_{12}S_{12}^{*}+S_{j_{5}}S_{j_{5}}^{*},$
$P_{e_{12}}=S_{g_{2}}^{*}S_{g_{2}}=S_{e_{1}}^{*}S_{e_{1}}=S_{j_{3}}^{*}S_{j_{3}}=S_{e_{2}}S_{e_{2}}^{*}+S_{g_{3}}S_{g_{3}}^{*}+S_{6}S_{6}^{*}+S_{i_{5}}S_{i_{5}}^{*}+S_{13}S_{13}^{*},$
and at this point, we are not going to have a detailed proof of the above relations, since it is easy to see them. This provide us with the desired Cuntz-Krieger $\mathcal{G}(\Pi_{3})$-family, and shows that the $C^{*}(S,P)$ is an infinite-dimensional $C^{*}$-algebra.
Now in order to proceed further, let us see what we have. For $n=2$, we have $S_{i}=\sum_{j=1}^{\infty}\prescript{i}{}{E}_{\mathcal{E}j-A,3j-D}$, and for $n=3$ case we have $S_{i}=\sum_{j=1}^{\infty}\prescript{i}{}{E}_{\mathcal{E}j-A,8j-D}$, and $S_{i}=\sum_{j=1}^{\infty}\prescript{i}{}{E}_{\mathcal{E}j-A,15j-D}$ for the $n=4$ case, and so on. So, we get a sequence of numbers $3,8,15,24,35,\cdots$. In order to have a defining rule for this sequence, we use the recurrence relations $a_{n-2}+2n=a_{n-1}$, and $b_{n-2}+1=b_{n-1}$, for $n\in\{2,\cdots\}$, together with conditional relations $a_{0}=1$ and $b_{0}=2$.
Now consider the recurrence relation $h_{n}+2n+1=h_{n+1}$, for $h_{n}=b_{n-2}+a_{n-2}$ and the conditional relation $h_{2}=b_{0}+a_{0}=3$.
Hence, we get the following relation
$S=\{S_{i}:=\sum_{j=1}^{\infty}\prescript{i}{}{E}_{\mathcal{E}j-A,h_{n}j-D}\mid\ \text{for fixed}\ 1\leq i\leq\frac{(n^{3}+n^{2})(n-1)}{2}\},$
for $h_{n}$ as above,and $n\in\{2,\cdots\}$. Now we can proceed by induction on $n$. For $n=2,3$, we already have seen that the assertion holds. Suppose that for $\mathcal{G}(\Pi_{i})$, $i\in\{2,3,\cdots\}$ (4.1.2) satisfies.
Now the claim is that for $\mathcal{G}(\Pi_{n+1})$ we have
$S=\{S_{i}:=\sum_{j=1}^{\infty}\prescript{i}{}{E}_{\mathcal{E}j-A,h_{n+1}j-D}\mid\ \text{for fixed}\ 1\leq i\leq\frac{(n^{3}+n^{2})(n-1)}{2}\},$
which is almost clear, since we have $(n+1)^{2}-1=n^{2}-1+2n+1=h_{n}+2n+1=h_{n+1},$ and we are done with the induction steps!
@@
In other case, it is easy to see that $10$-qubit in the quantum system introduced in this paper will be equivalent to the 1 MB=1024 Bytes of the classical systems.